# Quantum Machine Learning for Digital Health? A Systematic Review

### v1

@@ 6 Conclusion
Digital health aims to transform access, affordability and quality of healthcare. While classical machine learning methods in health are approaching commercialization, we find an exponentially growing number of studies advocating the use of QML in health. While assessing the future potential of use-case discovery in digital health is out of scope for a systematic review, we assess current-state evidence consisting of 4915 studies. We find most applications are focussed on clinical decision support and little attention is given to health service delivery and public health use-cases. Of eligible studies, we appraise study quality yielding 16 robust studies which analyze QML applications in realistic operating environments. However, we find that this synthesized empirical evidence does not establish clear trends in performance scaling of QML algorithms, and does not contextualize statistical fluctuations in classical vs. quantum analysis. Most QML algorithms were restricted to linear quantum models. Efficient data encoding remains elusive unless very specific hardware capabilities are assumed, while the impacts of data preprocessing on QML study design are unaddressed. Enabling meaningful use-case discovery for QML in digital health thus requires improved guidelines for empirical investigations, rationalizing the choice of QML structures on classical datasets, and establishing performance scaling under realistic operating conditions.
@@ Appendix A Quantum background
Quantum convolutional neural networks (QCNNs) can similarly be understood as a tunable quantum channel $\tilde{\eta}_{\theta}$ that is composed of many smaller channels $\eta_{\theta_{i}}$, where quantum channels are introduced in Appendix A. For each $i$-th tunable layer, the algorithm’s structure can be written as $\tilde{\eta}_{\theta}(\rho):=\bigcirc_{i}\eta_{\theta_{i}}(\rho)=\ldots\eta_{\theta_{3}}(\eta_{\theta_{2}}(\eta_{\theta_{1}}(\rho)))\ldots$. In typical formulations such as that in Fig. 1(d), in each $i$-th layer, we measure half of the remaining qubits in that layer, forcing these measured qubits to be reduced to classical bits. Consequently, these channels $\eta_{\theta_{i}}$ are defined on an increasingly smaller number of qubits as $i$ increases, until only one qubit is left. The channels for each layer $\eta_{\theta_{i}}$ are typically non-unitary, meaning that unlike quantum gates, these operations cannot be reversed or ‘undone’. An example of a non-unitary channel is where the measurement outcomes of a pooling layer dictate how gates are applied to the remaining qubits in the next layer [58]. If we want to extract $f(x,\theta)=\mathrm{Tr}\left[\tilde{\eta}_{\theta}(\rho_{x})\hat{O}\right]$, then in general it appears that tunable parameters and data-dependent operations cannot be separated for QCNNs. However, if mid-circuit QCNN measurements do not affect future quantum operations and can all be safely pushed to the end of the QCNN circuit, then averaging over mid-circuit measurements can be implemented entirely in classical post-processing and the layers $\eta_{\theta_{i}}$ can be unitary (i.e. $\eta_{\theta_{i}}(\rho):=V(\theta_{i})\rho V^{\dagger}(\theta_{i})$). In this regime, we recover a linear quantum model for QCNNs,
@@
Appendix B PRISMA Checklist
Appendix C Nomenclature
@@
Appendix D Search strategies
Our database selection captures publishing practices in healthcare and medicine but also in quantum computing. While PubMed and Embase are popularly used in health, we additionally include IEEE and Scopus which typically index computer science, physics and engineering journals. Physics-centric publications such as APS journals and Quantum are indexed in Scopus. Meanwhile we independently search quantum physics preprint server arXiv (quant-ph) to capture recent and unindexed quantum computing and quantum machine learning literature.
Appendix E Data extraction template

### v2

@@ 7 Conclusion
Digital health aims to transform access, affordability and quality of healthcare. As classical machine learning methods in health approach commercialization, we find an exponentially growing number of studies advocating the use of QML in health. Our work is the first systematic review that examines the strength of empirical evidence to support these claims using a database of 4915 studies. We find most applications are focussed on clinical decision support and comparatively little attention is given to health service delivery and public health use-cases. Of eligible studies, we appraise study quality yielding 16 robust studies which analyze QML applications in realistic operating environments. Despite this, we find that synthesized empirical evidence does not establish clear trends in performance benefits of QML algorithms over classical methods. Even leaving aside the issue of classical comparators, this synthesized evidence additionally does not establish scalability or robustness of QML performance. To this end, we propose minimal requirements for empirical studies for claiming empirical advantage for QML algorithms. We reiterate that enabling meaningful use-case discovery for QML in digital health requires new research to rationalize the choice of QML structures on classical datasets, define appropriate benchmarks, and establish performance scaling under realistic operating conditions. An update to the search period of our review enables us to systematically track changes in this evidence base in the near future.
@@ Appendix A Quantum background
Quantum convolutional neural networks (QCNNs) can similarly be understood as a tunable quantum channel $\tilde{\eta}_{\theta}$ that is composed of many smaller channels $\eta_{\theta_{i}}$, where quantum channels are introduced earlier. For each $i$-th tunable layer, the algorithm’s structure can be written as $\tilde{\eta}_{\theta}(\rho):=\bigcirc_{i}\eta_{\theta_{i}}(\rho)=\ldots\eta_{\theta_{3}}(\eta_{\theta_{2}}(\eta_{\theta_{1}}(\rho)))\ldots$. In typical formulations such as that in Fig. 1(d), in each $i$-th layer, we measure half of the remaining qubits in that layer, forcing these measured qubits to be reduced to classical bits. Consequently, these channels $\eta_{\theta_{i}}$ are defined on an increasingly smaller number of qubits as $i$ increases, until only one qubit is left. The channels for each layer $\eta_{\theta_{i}}$ are typically non-unitary, meaning that unlike quantum gates, these operations cannot be reversed or ‘undone’. An example of a non-unitary channel is where the measurement outcomes of a pooling layer dictate how gates are applied to the remaining qubits in the next layer [71]. If we want to extract $f(x,\theta)=\mathrm{Tr}\left[\tilde{\eta}_{\theta}(\rho_{x})\hat{O}\right]$, then in general it appears that tunable parameters and data-dependent operations cannot be separated for QCNNs. However, if mid-circuit QCNN measurements do not affect future quantum operations and can all be safely pushed to the end of the QCNN circuit, then averaging over mid-circuit measurements can be implemented entirely in classical post-processing and the layers $\eta_{\theta_{i}}$ can be unitary (i.e. $\eta_{\theta_{i}}(\rho):=V(\theta_{i})\rho V^{\dagger}(\theta_{i})$). In this regime, we recover a linear quantum model for QCNNs,
@@
Appendix B List of synthesized studies, code and datasets
For sixteen synthesized studies, we report available open-source code and datasets, where available with NR./UR. indicating Not Reported / Upon Request.
Synthesized Study Ref. Code Data
Nyugen 2018 [57] NR./UR. Synthetic
Piat 2018 [26] NR./UR. Private
Yano 2020 [51] NR./UR. UCI MLR Heart Disease Data.
UCI Wisconsin Breast Cancer Data.
Niraula 2021 [63] NR./UR. Private: RTOG0617 dataset collected under NCT00533949-D1 (c.f. NCTN)
Krunic 2022 [19] NR./UR. Private: Optum de-identified EHR dataset
Landman 2022 [52] NR./UR. MEDMINST Data.
Moradi 2022 [54] Git repo UCI Pediatric Bone Marrow transplant Data.
UCI Wisconsin Breast Cancer Data.
Kaggle Heart Failure Data.
Das 2023 [49] NR./UR. Private: MRIs of human ear collected in laboratory
Guddanti 2023 [25] Git repo Kaggle Pneumonia Chest X Ray Data
Kawaguchi 2023 [50] NR./UR. Kaggle Pima Indian Diabetes Data.
UCI MLR Heart Disease Data.
Moradi 2023 [55] Git repo Papp L et. al. Data. Grahovac M et. al. Data. Papp L, Pötsch N et al. Data.
Qu 2023 [48] Git repo MIT-BIH Arrhythmia & ST Change, European ST-T, Sudden Cardiac Death.
Asiwga 2024 [62] NR./UR. UCI Wisconsin Breast Cancer Data.
Cherrat 2024 [59] NR./UR. MEDMINST Data.
Kazdaghli 2024 [53] Git repo Synthetic, EHR MIMIC-III Data.
Choi 2024 [56] Git repo Private survey datasets.
Appendix C PRISMA Checklist
Appendix D Nomenclature
@@
Appendix E Search strategies
Our database selection captures publishing practices in healthcare/medicine but also in quantum computing. While PubMed and Embase are popular in health, we additionally include IEEE and Scopus which typically index computer science, physics and engineering journals. Physics-centric publications, e.g. APS journals and Quantum, are indexed in Scopus. We additionally search quantum physics preprint server arXiv (quant-ph) to capture recent and unindexed quantum computing and QML literature.
Appendix F Data extraction template